OpenStudy (anonymous):

The traffic light at the center of town is green for 45 seconds, then yellow for 5 seconds, then red for 40 seconds, at which point the cycle repeats. If a car approaches the traffic light at a random time, what is the probability that the traffic light will be green? A. 1/9 B.1/8 C.1/2 D.8/9 Please select the best answer from the choices provided A B C D im having a hard time!! help >.<

OpenStudy (anonymous):

i am thinking

OpenStudy (anonymous):

give me a medal if you want an answer

OpenStudy (anonymous):

i will! but please explain it for me too >.>

OpenStudy (anonymous):

add all the seconds and stuff

OpenStudy (anonymous):

mkay thats 90

OpenStudy (anonymous):

is it multiple answer

OpenStudy (anonymous):

uhm..no o_o

OpenStudy (anonymous):

then how long it the seconds of the green light ?

OpenStudy (anonymous):

45 >.>

OpenStudy (the_fizicx99):

Hm, the only formula I can think of is \(\ \sf \dfrac{Favorable~Outcome}{Total~Number~Of~Outcomes} \) So favorable outcome is a green light, 45 seconds, and the total possibility is having a green, yellow and red light! So that would be like having 45 + 5 + 90! :O Note: 90! Isn't a permutation :3 So that would be: \(\ \sf \dfrac{45}{ 45 + 5 + 40} \rightarrow \dfrac{45}{90} = ~ ? \)

OpenStudy (the_fizicx99):

And wot is 45/90? >.>

OpenStudy (anonymous):

2! sooo 1/2??

OpenStudy (anonymous):

and then reduce it as low as possible

OpenStudy (the_fizicx99):

o.0 How is 45/90 equal to 2? >.>

OpenStudy (anonymous):

you're all down now right?

OpenStudy (anonymous):

0.5 >.> <.<

OpenStudy (the_fizicx99):

Hun I think you meant the 2nd options :> Yesh it's 1/2 ;) 0.5 = 1/2 :>

OpenStudy (anonymous):

ugh ty so much! :)

OpenStudy (the_fizicx99):

Yw :>